# RefLoRA: Refactored Low-Rank Adaptation for Efficient Fine-Tuning of Large Models

### v1

@@ B.5 Proof of Theorem 4
It follows from (B.3) that $\mathbf{S}_{t}=\tilde{\mathbf{S}}_{t}$ is the unique global minima of this lower bound. ∎
@@ B.7 Useful facts
Let $\mathbf{X}=\mathbf{A}_{t}^{\top}\mathbf{A}_{t}$ and $\mathbf{Y}=\mathbf{B}_{t}^{\top}\mathbf{B}_{t}$ in (24) completes the proof. ∎
Theorem 8 (Descartes’ rule of signs [9]).
@@ Appendix C Algorithm pseudocodes
The Appendix provides the step-by-step pseudocodes for RefLoRA and its lightweight version RefLoRA-S. As our algorithms rely on Assumption 1 but LoRA intialize $\mathbf{B}_{0}=\mathbf{0}$ by default, one can either utilizes other full-rank initialization schemes [41, 34, 56], or warmup the algorithm for $t_{w}>0$ steps until $\mathbf{A}_{t}$ and $\mathbf{B}_{t}$ both satisfy full column rank. Without loss of generality, our pseudocodes assumes $\mathbf{A}_{0}$ and $\mathbf{B}_{0}$ are already full-rank.
Input: Loss $\ell$, pre-trained weight $\mathbf{W}^{\mathrm{pre}}$, maximum iterations $T$, learning rate $\eta$.
@@
4             Precondition gradients $\mathbf{G}_{A}=\nabla\ell(\mathbf{W}_{t})\mathbf{B}_{t}\tilde{\mathbf{S}}_{t}^{-1},~{}\mathbf{G}_{B}=\nabla\ell(\mathbf{W}_{t})^{\top}\mathbf{A}_{t}\tilde{\mathbf{S}}_{t}$ where $\mathbf{W}_{t}:=\mathbf{W}^{\mathrm{pre}}+\mathbf{A}_{t}\mathbf{B}_{t}^{\top}$;
@@
9             Update $\mathbf{A}_{t+1}=\tilde{\mathbf{A}}_{t}-\eta\nabla\ell(\mathbf{W}_{t})\tilde{\mathbf{B}}_{t},~{}\mathbf{B}_{t+1}=\tilde{\mathbf{B}}_{t}-\eta\nabla\ell(\mathbf{W}_{t})^{\top}\tilde{\mathbf{A}}_{t}$ where $\mathbf{W}_{t}:=\mathbf{W}^{\mathrm{pre}}+\mathbf{A}_{t}\mathbf{B}_{t}^{\top}$;
@@
Input: Loss $\ell$, pre-trained weight $\mathbf{W}^{\mathrm{pre}}$, maximum iterations $T$, learning rate $\eta$.
@@
5             Update $\mathbf{A}_{t+1}=\text{AdaptOpt}(\tilde{\mathbf{A}}_{t},\eta,\nabla\ell(\mathbf{W}_{t})\tilde{\mathbf{B}}_{t},t)$, $\mathbf{B}_{t+1}=\text{AdaptOpt}(\mathbf{B}_{t},\eta,\nabla\ell(\mathbf{W}_{t})^{\top}\tilde{\mathbf{A}}_{t},t)$ where $\mathbf{W}_{t}:=\mathbf{W}^{\mathrm{pre}}+\mathbf{A}_{t}\mathbf{B}_{t}^{\top}$;
@@
8             Update $\mathbf{A}_{t+1}=\tilde{\mathbf{A}}_{t}-\eta\nabla\ell(\mathbf{W}_{t})\tilde{\mathbf{B}}_{t},~{}\tilde{\mathbf{B}}_{t+1}=\tilde{\mathbf{B}}_{t}-\eta\nabla\ell(\mathbf{W}_{t})^{\top}\tilde{\mathbf{A}}_{t}$ where $\mathbf{W}_{t}:=\mathbf{W}^{\mathrm{pre}}+\mathbf{A}_{t}\mathbf{B}_{t}^{\top}$;
@@ D.1 Platforms
All the experiments are conducted on a desktop equipped with an NVIDIA RTX A5000 GPU, and a server with NVIDIA A40 and A100 GPUs. The codes for synthetical tests are written with MATLAB, and codes for LLM-related experiments are in PyTorch. In addition, our implementation of LLM tests are based on [23, 66].
@@ D.2 Setups for visualization in Figure 1
where $\mathbf{X}\in^{n\times k}$ and $\mathbf{Y}\in^{m\times k}$ are random data matrices with entries generated from standard Gaussian distribution $\mathcal{N}(0,1)$. The corresponding LoRA objective is
$\min_{\mathbf{A},\mathbf{B}}\|\mathbf{Y}-(\mathbf{W}^{\mathrm{pre}}+\mathbf{A}\mathbf{B}^{\top})\mathbf{X}\|_{\mathrm{F}}^{2}.$ (25)
For simplicity, we set $m=n=k=2$, $r=1$, and $\mathbf{W}^{\mathrm{pre}}=\mathbf{0}$. $\mathbf{A}_{0}$ and $\mathbf{B}_{0}$ are randomly initialized from $\mathcal{N}(0,10)$ and $\mathcal{N}(0,\frac{1}{10})$.
Although it is well-known that linear regression has a closed-form solution through least squares, we consider it here because its Lipschitz smoothness constant can be computed analytically as $L=\|\mathbf{X}\mathbf{X}^{\top}\|_{2}$, allowing us to track the quadratic and alternative upper bounds.
@@ D.3 Setups for matrix factorization
where $\mathbf{Y}\in^{m\times n}$ is a given low-rank matrix. It can be viewed as a special case of the linear model (25) with $k=n$, $\mathbf{X}=\mathbf{I}_{n}$, and $\mathbf{W}^{\mathrm{pre}}=\mathbf{0}$. This test utilizes $m=128$, $n=100$, and $r=8$. The low-rank matrix $\mathbf{Y}$ is generated from standard Gaussian $\mathcal{N}(0,1)$, and then truncated to the largest $r$ singular values. Following the standard LoRA initialization, $\mathbf{A}_{0}$ is sampled from standard Gaussian $\mathcal{N}(0,1)$, and $\mathbf{B}_{0}=\mathbf{0}$. Standard GD is employed in LoRA for $\forall t$, while RefLoRA and ScaledGD are applied when $t>0$.
@@ D.4 Details of Datasets
GLUE benchmark. The General Language Understanding Evaluation (GLUE) benchmark is designed to provide a general-purpose evaluation of natural language understanding (NLU) [55]. Those adopted in our work include
  • MNLI [59] (Multi-Genre Natural Language Inference) tests a model’s ability to perform natural language inference across different genres of text.
  • SST-2 [49] (Stanford Sentiment Treebank) is a sentiment analysis dataset with binary labels.
  • MRPC [11] (Microsoft Research Paraphrase Corpus) focuses on paraphrase detection; i.e. identifying whether two sentences are semantically equivalent.
  • CoLA [58] (Corpus of Linguistic Acceptability) requires models to judge whether a sentence is linguistically acceptable.
  • QNLI [43] (Question Natural Language Inference) is a question-answering dataset converted to a binary inference task.
@@
  • STS-B [6] (Semantic Textual Similarity Benchmark) evaluates the textual similarity of sentence pairs on a continuous scale.
@@
MNLI Natural language inference 393k 20k Accuracy
SST-2 Sentiment analysis 67k 1.8k Matched & mismatched accuracies
@@
  • WG [46] (WinoGrande) is a challenging dataset designed to reduce annotation artifacts present in traditional Winograd schemas.
  • PIQA [4] (Physical Interaction QA) assesses knowledge of physical commonsense and intuitive physics.
  • SIQA [47] (SOCIAL-I-QA) focuses on social interaction and social commonsense reasoning.
  • HS [64] (HellaSwag) aims at evaluating grounded commonsense inference for multiple-choice sentence completion.
  • ARC [8] (AI2 Reasoning Challenge) contains grade-school level science questions, split into ARCe (ARC-easy) and ARCc (ARC-challenge), based on difficulty.
  • OpenbookQA [42] involves multiple-choice science questions that require integrating commonsense and scientific facts.
@@ D.5 Details on LLMs
DeBERTaV3-base [20] is a transformer-based language model with 184 million parameters. The model checkpoint [^4] https://huggingface.co/microsoft/deberta-v3-base is released under the MIT license.
@@
LLaMA-7B [53] is a decoder-only transformer model, which is part of the LLaMA (Large Language Model Meta AI) series. The chekpoint [^6] https://huggingface.co/huggyllama/llama-7b is intended for research use under a non-commercial license.
LLaMA2-7B [54] is a refined successor to LLaMA. Its checkpoint [^7] https://huggingface.co/meta-llama/Llama-2-7b is under a permissive license for both research and commercial use.
LLaMA3-8B [16] is part of the third generation of LLaMA series. The checkpoint [^8] https://huggingface.co/meta-llama/Meta-Llama-3-8B is released under a permissive Meta license for both research and commercial applications.
@@ D.6 Hyperparameters for fine-tuning LLMs
GLUE setup follows from [22, 66], where LoRA is applied to all linear modules. The results for Full FT, BitFit, Adapters, LoRA, DoRA and AdaLoRA in Table 2 are taken from  [66], while the remaining results are obtained from our experiments. We fix the LoRA rank as $r=8$ and scaling factor as $\alpha=8$, and search the optimal learning rate from $\eta\in\{1\times 10^{-3},8\times 10^{-4},4\times 10^{-4}\}$. Batch size is fixed as $32$ for all datasets. The default AdamW [40] optimizer and linear learning rate scheduler are utilized for all tests. Other hyperparameters are gathered in Table 6. Note that RefLoRA requires less fine-tuning epochs compared other LoRA variants due to its fast convergence.
@@
Commonsense reasoning setup is from [23, 61], where LoRA is attached to linear projections in transformers’ self-attention and feedforward modules. The results for ChatGPT, LoRA, and DoRA in Table 3 are taken from  [61], while the remaining are acquired through our tests. We test with LoRA ranks $r\in\{16,32\}$ and scaling factor fixed as $\alpha=2r$. The learning rate is tuned from $\eta\in\{8\times 10^{-5},1\times 10^{-4},2\times 10^{-4},3\times 10^{-4}\}$. The tuned learning rate per model can be found in Table 7. The number of fine-tuning epochs is set to $2$, and batch size is $16$ for all tests. The default AdamW [40] optimizer and linear learning rate scheduler are utilized with $100$ warmup steps. Dropout rate is $0.05$. The remaining hyperparameters are set to the default values used in [23].
We also attempted to include results for LoRA-Pro [57] and LoRA-RITE [63]; however, the former incurred runtime costs that exceeded the limits of our resources; and the latter struggles on convergence even if we have tuned the learning rates from $10^{-6}$ to $10^{-3}$. For these reasons, we omitted them from our final results.

### v2

@@ B.5 Proof of Theorem 4
It follows from (B.3) that $\mathbf{S}_{t}=\tilde{\mathbf{S}}_{t}$ is the unique global minimum of this lower bound. ∎
@@ B.7 Useful facts
Letting $\mathbf{X}=\mathbf{A}_{t}^{\top}\mathbf{A}_{t}$ and $\mathbf{Y}=\mathbf{B}_{t}^{\top}\mathbf{B}_{t}$ in (24) completes the proof. ∎
Theorem 8 (Descartes’ rule of signs Descartes ).
@@ Appendix C Algorithm pseudocodes
The Appendix provides the step-by-step pseudocodes for RefLoRA and its lightweight version RefLoRA-S. As our algorithms rely on Assumption 1 but LoRA intialize $\mathbf{B}_{0}=\mathbf{0}$ by default, one can either utilizes other full-rank initialization schemes PiSSA ; LoftQ ; LoRA-GA , or warm up the algorithm for $t_{w}>0$ steps until $\mathbf{A}_{t}$ and $\mathbf{B}_{t}$ both satisfy full column rank. Without loss of generality, our pseudocodes assume $\mathbf{A}_{0}$ and $\mathbf{B}_{0}$ are already full-rank.
Input: Loss $\ell$, pre-trained weight $\mathbf{W}^{\mathrm{pt}}$, maximum iterations $T$, and learning rate $\eta$.
@@
4      Precondition gradients $\mathbf{G}_{A}=\nabla\ell(\mathbf{W}_{t})\mathbf{B}_{t}\tilde{\mathbf{S}}_{t}^{-1},~\mathbf{G}_{B}=\nabla\ell(\mathbf{W}_{t})^{\top}\mathbf{A}_{t}\tilde{\mathbf{S}}_{t}$ where $\mathbf{W}_{t}:=\mathbf{W}^{\mathrm{pt}}+\mathbf{A}_{t}\mathbf{B}_{t}^{\top}$;
@@
9      Update $\mathbf{A}_{t+1}=\tilde{\mathbf{A}}_{t}-\eta\nabla\ell(\mathbf{W}_{t})\tilde{\mathbf{B}}_{t},~\mathbf{B}_{t+1}=\tilde{\mathbf{B}}_{t}-\eta\nabla\ell(\mathbf{W}_{t})^{\top}\tilde{\mathbf{A}}_{t}$ where $\mathbf{W}_{t}:=\mathbf{W}^{\mathrm{pt}}+\mathbf{A}_{t}\mathbf{B}_{t}^{\top}$;
@@
Input: Loss $\ell$, pre-trained weight $\mathbf{W}^{\mathrm{pt}}$, maximum iterations $T$, and learning rate $\eta$.
@@
5      Update $\mathbf{A}_{t+1}=\text{AdaptOpt}(\tilde{\mathbf{A}}_{t},\eta,\nabla\ell(\mathbf{W}_{t})\tilde{\mathbf{B}}_{t},t)$, $\mathbf{B}_{t+1}=\text{AdaptOpt}(\mathbf{B}_{t},\eta,\nabla\ell(\mathbf{W}_{t})^{\top}\tilde{\mathbf{A}}_{t},t)$ where $\mathbf{W}_{t}:=\mathbf{W}^{\mathrm{pt}}+\mathbf{A}_{t}\mathbf{B}_{t}^{\top}$;
@@
8      Update $\mathbf{A}_{t+1}=\tilde{\mathbf{A}}_{t}-\eta\nabla\ell(\mathbf{W}_{t})\tilde{\mathbf{B}}_{t},~\tilde{\mathbf{B}}_{t+1}=\tilde{\mathbf{B}}_{t}-\eta\nabla\ell(\mathbf{W}_{t})^{\top}\tilde{\mathbf{A}}_{t}$ where $\mathbf{W}_{t}:=\mathbf{W}^{\mathrm{pt}}+\mathbf{A}_{t}\mathbf{B}_{t}^{\top}$;
@@ D.1 Platforms
All the experiments are conducted on a desktop equipped with an NVIDIA RTX A5000 GPU, and a server with NVIDIA A40 and A100 GPUs. The codes for synthetical tests are written with MATLAB, and codes for LLM-related experiments are in PyTorch. In addition, our implementation of LLM tests are based on hu2023llm ; AdaLoRA .
@@ D.2 Setups for visualization in Figure 1
where $\mathbf{X}\in\mathbb{R}^{n\times k}$ and $\mathbf{Y}\in\mathbb{R}^{m\times k}$ are random data matrices with entries generated from standard Gaussian distribution $\mathcal{N}(0,1)$. The corresponding LoRA objective is
$\min_{\mathbf{A},\mathbf{B}}\|\mathbf{Y}-(\mathbf{W}^{\mathrm{pt}}+\mathbf{A}\mathbf{B}^{\top})\mathbf{X}\|_{\mathrm{F}}^{2}.$ (25)
For simplicity, we set $m=n=k=2$, $r=1$, and $\mathbf{W}^{\mathrm{pt}}=\mathbf{0}$. $\mathbf{A}_{0}$ and $\mathbf{B}_{0}$ are randomly initialized from $\mathcal{N}(0,10)$ and $\mathcal{N}(0,\frac{1}{10})$.
Although it is well-known that linear regression has a closed-form solution through least squares, we consider it here because its Lipschitz smoothness constant can be computed analytically as $L=\|\mathbf{X}\mathbf{X}^{\top}\|_{2}$, allowing us to track the upper bound (7).
@@ D.3 Setups for matrix factorization
where $\mathbf{Y}\in\mathbb{R}^{m\times n}$ is a given low-rank matrix. It can be viewed as a special case of the linear model (25) with $k=n$, $\mathbf{X}=\mathbf{I}_{n}$, and $\mathbf{W}^{\mathrm{pt}}=\mathbf{0}$. This test utilizes $m=128$, $n=100$, and $r=8$. The low-rank matrix $\mathbf{Y}$ is generated from standard Gaussian $\mathcal{N}(0,1)$, and then truncated to the largest $r$ singular values. Following the standard LoRA initialization, $\mathbf{A}_{0}$ is sampled from standard Gaussian $\mathcal{N}(0,1)$, and $\mathbf{B}_{0}=\mathbf{0}$. Standard GD is employed in LoRA for $\forall t$, while RefLoRA and ScaledGD are applied when $t>0$.
@@ D.4 Details of Datasets
GLUE benchmark. The General Language Understanding Evaluation (GLUE) benchmark is designed to provide a general-purpose evaluation of natural language understanding (NLU) wang2018glue . Those adopted in our work include
  • MNLI mnli (Multi-Genre Natural Language Inference) tests a model’s ability to perform natural language inference across different genres of text.
  • SST-2 sst2 (Stanford Sentiment Treebank) is a sentiment analysis dataset with binary labels.
  • MRPC mrpc (Microsoft Research Paraphrase Corpus) focuses on paraphrase detection; i.e. identifying whether two sentences are semantically equivalent.
  • CoLA cola (Corpus of Linguistic Acceptability) requires models to judge whether a sentence is linguistically acceptable.
  • QNLI qnli (Question Natural Language Inference) is a question-answering dataset converted to a binary inference task.
@@
  • STS-B stsb (Semantic Textual Similarity Benchmark) evaluates the textual similarity of sentence pairs on a continuous scale.
@@
MNLI Natural language inference 393k 20k Matched & mismatched accuracies
SST-2 Sentiment analysis 67k 1.8k Accuracy
@@
  • BoolQ boolq (Boolean Questions) is a reading comprehension dataset of yes/no questions paired with Wikipedia passages, testing a model’s ability to extract and reason over text.
  • WG winogrande (WinoGrande) is a challenging dataset designed to reduce annotation artifacts present in traditional Winograd schemas.
  • PIQA piqa (Physical Interaction QA) assesses knowledge of physical commonsense and intuitive physics.
  • SIQA siqa (SOCIAL-I-QA) focuses on social interaction and social commonsense reasoning.
  • HS hellaswag (HellaSwag) aims at evaluating grounded commonsense inference for multiple-choice sentence completion.
  • ARC arc (AI2 Reasoning Challenge) contains grade-school level science questions, split into ARCe (ARC-easy) and ARCc (ARC-challenge), based on difficulty.
  • OpenbookQA obqa involves multiple-choice science questions that require integrating commonsense and scientific facts.
@@ D.5 Details on LLMs
DeBERTaV3-base DeBERTaV3 is a transformer-based language model with 184 million parameters. The model checkpoint [^4] https://huggingface.co/microsoft/deberta-v3-base is released under the MIT license.
@@
LLaMA-7B llama is a decoder-only transformer model, which is part of the LLaMA (Large Language Model Meta AI) series. The chekpoint [^6] https://huggingface.co/huggyllama/llama-7b is intended for research use under a non-commercial license.
LLaMA2-7B llama2 is a refined successor to LLaMA. Its checkpoint [^7] https://huggingface.co/meta-llama/Llama-2-7b is under a permissive license for both research and commercial use.
LLaMA3-8B Llama3 is part of the third generation of LLaMA series. The checkpoint [^8] https://huggingface.co/meta-llama/Meta-Llama-3-8B is released under a permissive Meta license for both research and commercial applications.
Stable Diffusion V1.4 stable-diffusion is a latent text-to-image diffusion model released by CompVis, Stability AI, and Runway. The checkpoint [^9] https://huggingface.co/CompVis/stable-diffusion-v1-4 is made available under the CreativeML-OpenRAIL-M license.
@@ D.6 Hyperparameters for fine-tuning LLMs
GLUE setup follows from LoRA ; AdaLoRA , where LoRA is applied to all linear modules. The results for Full FT, BitFit, Adapters, LoRA, DoRA and AdaLoRA in Table 2 are taken from  AdaLoRA , while the remaining results are obtained from our experiments. We fix the LoRA rank as $r=8$ and scaling factor as $\alpha=8$, and search the optimal learning rate from $\eta\in\{1\times 10^{-3},8\times 10^{-4},4\times 10^{-4}\}$. Batch size is fixed as $32$ for all datasets. The default AdamW AdamW optimizer and linear learning rate scheduler are utilized for all tests. Other hyperparameters are gathered in Table 7. Note that RefLoRA requires less fine-tuning epochs compared other LoRA variants due to its fast convergence.
@@
Commonsense reasoning setup is from hu2023llm ; DoRA , where LoRA is attached to linear projections in transformers’ self-attention and feedforward modules. The results for ChatGPT, LoRA, and DoRA in Table 3 are taken from  DoRA , while the remaining are acquired through our tests. We test with LoRA ranks $r\in\{16,32\}$ and scaling factor fixed as $\alpha=2r$. The learning rate is tuned from $\eta\in\{8\times 10^{-5},1\times 10^{-4},2\times 10^{-4},3\times 10^{-4}\}$. The tuned learning rate per model can be found in Table 8. The number of fine-tuning epochs is set to $2$, and batch size is $16$ for all tests. The default AdamW AdamW optimizer and linear learning rate scheduler are utilized with $100$ warmup steps. Dropout rate is $0.05$. The remaining hyperparameters are set to the default values used in hu2023llm .